If there are initially 1800 bacteria in a culture and the number of bacteria triple each hour, the number of bacteria after t hours can be found using. Y=1800(3)^t
1
Expert's answer
2019-08-29T09:25:27-0400

If we need to find the number of hours needed to grow to some exact number y, then:

3^t = Y/1800,

log3^t = log(Y/1800),

t×log3 = log(Y/1800),

t = log(Y/1800)/log3.